# Fractional Sobolev paths on Wasserstein spaces and their energy-minimizing particle representations

### v3

@@ 3.5. Characterization of geodesics via Besov regularity
See e.g. [2, Lemma 9.11]. The goal of this section is to present a similar characterization using the Besov regularity $|\cdot|_{b^{\alpha,p}}$, which subsequently allows us to give a dynamic reformulation of the Wasserstein metric in the next section. We begin with a simple observation, which will be proven later with the main lemma.
@@
Let $(\mathcal{X},d)$ be a metric space. Let $1\leq p<\infty$ and $10<\alpha<$ 1. For any (possibly discontinuous) $\gamma:[0,1]\to\mathcal{X}$ with $|\gamma|_{b^{\alpha,p}}<+\infty$, we have
@@
Let $(\mathcal{X},d)$ be a metric space. Let $1<p<\infty$ and $10<\alpha<$ 1. Then the following are equivalent:
@@ 3.6. Dynamic formulation of Wasserstein distance via Besov energy
As a consequence of Lemmas 3.11 and 3.12, we obtain a generalization of the metric Benamou–Brenier formula to the fractional setting. Its proof is similar to that for $W^{1,p}$-energy (e.g. [2, Theorem 9.13]).
@@
Let $(\mathcal{X},d)$ be a complete, separable, and geodesic metric space. Let $1<p<\infty$ and $10<\alpha<$ 1. Then for every $\mu,\nu\in P_{p}(\mathcal{X})$, we have
@@ 4. Counterexamples
This elementary example is well-known in stochastic analysis, see e.g. [20, Chapter I. Exercise (2.12)], and it is studied in detail in [1, Example 1.1]. Consider the following collection of measures on $\mathcal{X}\coloneqq\mathbb{R}$:
@@
Unlike the case $p=1$, where the 1-variation of $1$-Wasserstein curves can be realized by (optimal) lifts [1], this is not the case for $p$-variation of $p$-Wasserstein curves when $p>1$, even though in both cases, the norm is obtained by taking supremum over all partitions. The failure in the case $p>1$ arises because higher-order variations of different curves in the underlying space can potentially be achieved by different sequences of partitions, in contrast with $p=1$, where the $1$-variation of curves can be achieved by the same sequence of partitions with the shrinking mesh size. Recall (2.18), where $p=1$ is the only case that variation and its infinitesimal characterization coincide and capture the regularity “locally”. Here, we observe that higher-order variations behave more like Hölder regularity and the modulus of continuity, rather than 1-variation.
@@
$\mu_{t}\coloneqq\frac{1}{2}\delta_{\gamma_{t}^{1}}+\frac{1}{2}\delta_{\gamma_{t}^{2}},\qquad\forall t\in I\coloneqq[0,1],$ (4.4)
@@
Obviously, two-dimensional marginals of $\pi$ give the optimal coupling at all pairs of times with respect to all $W_{p}$, $p\geq 1$, thus $(\mu_{t})$ is compatible. By [1, Theorem 3.3] or direct computation, the $1$-variation of $(\mu_{t})$ with respect to $1$-Wasserstein metric satisfies:
@@
Specifically, here we construct a $\upgamma$-Hölder compatible curve $(\mu_{t})$ in $p$-Wasserstein space ($11/p<\upgamma<$ 1) by superimposing absolutely continuous curves in the underlying space. Here, $(\mu_{t})$ can also be seen as the solution to the continuity equation. For $(\mu_{t})$, there will be only one lift $\pi$ on the space of continuous paths, so any construction should lead to this lift. Since $(\mu_{t})$ forms a compatible collection, the lift here also enjoys the property that all of its two-dimensional marginals are optimal. We show that $\frac{1}{\upgamma}$-variation of $(\mu_{t})$ along the sequence of dyadic partitions has a non-zero limit, whereas this variation is zero for all curves in the underlying space (because they are of bounded 1-variation). From this, we will conclude that
@@
[…]
@@
Claim 3 ($\int|\gamma|^{p}_{W^{\alpha,p}}{\mathrm{\,d}}\pi<\infty$ for all $\alpha\in(\frac{1}{p},\upgamma)$). Here, we aim to show that, in contrast to the computed $W^{1,p}$-energy above, the $W^{\alpha,p}$-energy of the lift $\pi$ is finite. Let’s begin by calculating $|\gamma^{j}|_{b^{\alpha,p}}$. Note that each $t\mapsto\gamma_{t}^{j}$ is a piece-wise geodesic curve, for which we have already computed the $b^{\alpha,p}$-norm in Lemma 2.31. Using this result with $n=j+1$ and noticing that the only non-zero summand in the first sum in (2.57) is for $m=n$, we obtain
[…]
@@
Given the optimality of $\pi$ and (4.22), one can argue, by Lisini’s characterization of absolutely continuous curves [14], and the recent characterization of BV-curves in $1$-Wasserstein space [1, Corollary 4.1], that $t\mapsto\mu_{t}$ is not an absolutely continuous curve in $p$-Wasserstein space, $p>1$, and it is not a BV-curve in 1-Wasserstein space, respectively.
@@
Denote by $\pi$ and $\pi^{j}$ the only existing lifts of $(\mu_{t})$ and $(\mu_{t}^{j})$ on the space of continuous paths, respectively. The main observation here is that, while this splitting procedure doesn’t change the energy of the lifts (with respect to any norm), it may change the regularity of Wasserstein curves with respect to certain norms. Let us first start with an observation on compatibility.
@@
Figure 4. Example 4.4 yields non-compatible piece-wise geodesic curves $(\mu_{t}^{j})\in C([0,1];P_{p}(\mathbb{S}^{1}))$ indexed by $j$ by modifying a constant-speed geodesic $(\mu_{t})$ in the $p$-Wasserstein space on $\mathbb{S}^{1}$, where $p\in[1,\infty)$. Top: Initial measures defined in (4.35) and (4.36). Each $\mu_{t}^{j}$ consists of $2^{j+1}$ particles with equal mass arranged equidistantly on the circle with a perimeter equal to 2. As time runs, the particles rotate around the circle with constant speed 1. Bottom: Wasserstein distance from the initial measure as a function of time. The metric speed of the Wasserstein curves is 1 for a.e. $t\in[0,1]$ for all $j$, whereas the fractional Sobolev norm tends to 00 as $j\to\infty$.
@@
Observation 3 ($p$-variation of $(\mu_{t}^{j})$ decreases in $j$ for all $p>1$ while $1$-variation remains constant). The $1$-variation of the curve $(\mu_{t}^{j})$ with respect to $1$-Wasserstein distance is equal to 1 for all $j$. Also recalling [1, Theorem 3.3], we write
$11=|\mu^{j}|_{1\textrm{-}\mathrm{var}}=\int|\gamma|_{1\textrm{-}\mathrm{var}}{\mathrm{\,d}}\pi^{j}=\int|\gamma|_{1\textrm{-}\mathrm{var}}{\mathrm{\,d}}\pi=|\mu|_{1\textrm{-}\mathrm{var}}=$ 1 . (4.39)
Now take $p>1$ and note that the $p$-variation of the curve $(\mu_{t}^{j})$ with respect to $p$-Wasserstein metric is achieved by the partition $(\frac{i}{2^{j+1}})_{i\in\{0,\cdots 2^{j+1}\}}$. This can be verified using the interpolation result [10, Proposition 5.5]. We thus obtain
@@
Observation 4 ($W^{1,p}$-regularity of $(\mu_{t}^{j})$ remains constant). Consider $1<p<\infty$. Each $(\mu_{t}^{j})$ is an absolutely continuous curve in $p$-Wasserstein space with its metric speed $|\dot{\mu}_{t}^{j}|=1$ for almost all time for all $j$. Recalling [14, Theorem 5], we have here:
$11=|\mu^{j}|^{p}_{W^{1,p}}=\int|\gamma|^{p}_{W^{1,p}}{\mathrm{\,d}}\pi^{j}=\int|\gamma|^{p}_{W^{1,p}}{\mathrm{\,d}}\pi=|\mu|^{p}_{W^{1,p}}=$ 1 . (4.41)
Observation 5 ($W^{\alpha,p}$-regularity of $(\mu_{t}^{j})$ decreases in $j$). Consider $1<p<\infty$ and $11/p<\alpha<$ 1. Then
@@
This is the main counterexample of the paper, whose goal is to show that if we are given $(\mu_{t})\in W^{\alpha,p}(I;P_{p}(\mathcal{X}))$ with $11/p<\alpha<$ 1 but without the compatibility property, there can be no lift with finite $W^{\alpha,p}$-energy. Specifically, we construct a non-compatible curve $(\mu_{t})_{t\in[0,1]}$ in a $p$-Wasserstein space with finite $W^{\alpha,p}$-norm whose only existing lift $\pi$ has infinite $W^{\alpha,p}$-energy:
@@
The construction follows by combining Examples 4.4 and 4.3, leading to an example on a cylinder. Due to the similarities, we shall not repeat all the details. Here, by choosing $\upgamma=\alpha$ for the mass $w_{j}$ in (4.14), we let the $b^{\alpha,p}$-energy (4.24) of the lift $\pi$ diverge. Meanwhile, by modifying the measures with the particle splitting procedure explained in Example 4.4, we let the $b^{\alpha,p}$-norm of $(\mu_{t})$ decrease and eventually converge to a finite quantity. This procedure clearly won’t change the energy. In particular, the computations in the aforementioned examples can be used.
@@
[…]
@@
$\int_{\Gamma}|\gamma|^{p}_{W^{1,p}}{\mathrm{\,d}}\pi(\gamma)$ $=\int_{\Gamma}\int_{0}^{1}|\dot{\gamma}_{t}|^{p}{\mathrm{\,d}}t{\mathrm{\,d}}\pi(\gamma)=\bar{w}2^{p}\sum_{j=0}^{\infty}2^{j(p-\alpha p)}=+\infty.$ (4.47)
@@
$\int_{\Gamma}|\gamma|^{p}_{b^{\alpha,p}}{\mathrm{\,d}}\pi(\gamma)=\frac{\bar{w}}{2^{-\alpha p}-2^{-p}}\sum_{j=0}^{\infty}2^{j(\alpha p-\alpha p)}=+\infty.$ (4.48)
@@
[…]

### v4

@@ 3.5. Characterization of geodesics via Besov regularity
See e.g. [4, Lemma 9.11]. The goal of this section is to present a similar characterization using the Besov regularity $|\cdot|_{b^{\alpha,p}}$, which subsequently allows us to give a dynamic reformulation of the Wasserstein metric in the next section. We begin with a simple observation, which will be proven later with the main lemma.
@@
Let $(\mathcal{X},d)$ be a metric space. Let $1\leq p<\infty$ and $0<\alpha<1$. For any (possibly discontinuous) $\gamma:[0,1]\to\mathcal{X}$ with $|\gamma|_{b^{\alpha,p}}<+\infty$, we have
@@
Let $(\mathcal{X},d)$ be a metric space. Let $1<p<\infty$ and $0<\alpha<1$. Then the following are equivalent:
@@ 3.6. Dynamic formulation of Wasserstein distance via Besov energy
As a consequence of Lemmas 3.11 and 3.12, we obtain a generalization of the metric Benamou–Brenier formula to the fractional setting. Its proof is similar to that for $W^{1,p}$-energy (e.g. [4, Theorem 9.13]).
@@
Let $(\mathcal{X},d)$ be a complete, separable, and geodesic metric space. Let $1<p<\infty$ and $0<\alpha<1$. Then for every $\mu,\nu\in P_{p}(\mathcal{X})$, we have
@@ 4. Counterexamples
This elementary example is well-known in stochastic analysis, see e.g. [24, Chapter I. Exercise (2.12)], and it is studied in detail in [2, Example 1.1]. Consider the following collection of measures on $\mathcal{X}\coloneqq\mathbb{R}$:
@@
Unlike the case $p=1$, where the 1-variation of $1$-Wasserstein curves can be realized by (optimal) lifts [2], this is not the case for $p$-variation of $p$-Wasserstein curves when $p>1$, even though in both cases, the norm is obtained by taking supremum over all partitions. The failure in the case $p>1$ arises because higher-order variations of different curves in the underlying space can potentially be achieved by different sequences of partitions, in contrast with $p=1$, where the $1$-variation of curves can be achieved by the same sequence of partitions with the shrinking mesh size. Recall (2.18), where $p=1$ is the only case that variation and its infinitesimal characterization coincide and capture the regularity ``locally''. Here, we observe that higher-order variations behave more like Hölder regularity and the modulus of continuity, rather than 1-variation.
@@
$\mu_{t}\coloneqq\frac{1}{2}\delta_{\gamma_{t}^{1}}+\frac{1}{2}\delta_{\gamma_{t}^{2}},\qquad\forall t\in I\coloneqq[0,1],$
@@
Obviously, two-dimensional marginals of $\pi$ give the optimal coupling at all pairs of times with respect to all $W_{p}$, $p\geq 1$, thus $(\mu_{t})$ is compatible. By [2, Theorem 3.3] or direct computation, the $1$-variation of $(\mu_{t})$ with respect to $1$-Wasserstein metric satisfies:
@@
Specifically, here we construct a $\upgamma$-Hölder compatible curve $(\mu_{t})$ in $p$-Wasserstein space ($1/p<\upgamma<1$) by superimposing absolutely continuous curves in the underlying space. Here, $(\mu_{t})$ can also be seen as the solution to the continuity equation. For $(\mu_{t})$, there will be only one lift $\pi$ on the space of continuous paths, so any construction should lead to this lift. Since $(\mu_{t})$ forms a compatible collection, the lift here also enjoys the property that all of its two-dimensional marginals are optimal. We show that $\frac{1}{\upgamma}$-variation of $(\mu_{t})$ along the sequence of dyadic partitions has a non-zero limit, whereas this variation is zero for all curves in the underlying space (because they are of bounded 1-variation). From this, we will conclude that
@@
$w_{j}\coloneqq\bar{w}\,2^{-jp\upgamma},\qquad\bar{w}(p,\upgamma)\coloneqq 1-2^{-p\upgamma},$ (4.13)
@@
Claim 3 ($\int|\gamma|^{p}_{W^{\alpha,p}}{\mathrm{\,d}}\pi<\infty$ for all $\alpha\in(\frac{1}{p},\upgamma)$). Here, we aim to show that, in contrast to the computed $W^{1,p}$-energy above, the $W^{\alpha,p}$-energy of the lift $\pi$ is finite. Let's begin by calculating $|\gamma^{j}|_{b^{\alpha,p}}$. Note that each $t\mapsto\gamma_{t}^{j}$ is a piece-wise geodesic curve, for which we have already computed the $b^{\alpha,p}$-norm in Lemma 2.31. Using this result with $n=j+1$ and noticing that the only non-zero summand in the first sum in (2.55) is for $m=n$, we obtain
$|\gamma^{j}|_{b^{\alpha,p}}^{p}\,=\,2^{(j+1)(\alpha p)}+\frac{2^{(j+1)(\alpha p-1)}}{2^{(p-\alpha p)}-1}\sum_{i=0}^{2^{j+1}-1}1^{p}\,=\,\frac{1}{2^{-\alpha p}-2^{-p}}\,2^{j\alpha p}.$ (4.22)
@@
Given the optimality of $\pi$ and (4.21), one can argue, by Lisini's characterization of absolutely continuous curves [17], and the recent characterization of BV-curves in $1$-Wasserstein space [2, Corollary 4.1], that $t\mapsto\mu_{t}$ is not an absolutely continuous curve in $p$-Wasserstein space, $p>1$, and it is not a BV-curve in 1-Wasserstein space, respectively.
@@
Denote by $\pi$ and $\pi^{j}$ the only existing lifts of $(\mu_{t})$ and $(\mu_{t}^{j})$ on the space of continuous paths, respectively. The main observation here is that, while this splitting procedure doesn't change the energy of the lifts (with respect to any norm), it may change the regularity of Wasserstein curves with respect to certain norms. Let us first start with an observation on compatibility.
@@
Figure 4. Example 4.4 yields non-compatible piece-wise geodesic curves $(\mu_{t}^{j})\in C([0,1];P_{p}(\mathbb{S}^{1}))$ indexed by $j$ by modifying a constant-speed geodesic $(\mu_{t})$ in the $p$-Wasserstein space on $\mathbb{S}^{1}$, where $p\in[1,\infty)$. Top: Initial measures defined in (4.34) and (4.35). Each $\mu_{t}^{j}$ consists of $2^{j+1}$ particles with equal mass arranged equidistantly on the circle with a perimeter equal to 2. As time runs, the particles rotate around the circle with constant speed 1. Bottom: Wasserstein distance from the initial measure as a function of time. The metric speed of the Wasserstein curves is 1 for a.e. $t\in[0,1]$ for all $j$, whereas the fractional Sobolev norm tends to 0 as $j\to\infty$.
@@
Observation 3 ($p$-variation of $(\mu_{t}^{j})$ decreases in $j$ for all $p>1$ while $1$-variation remains constant). The $1$-variation of the curve $(\mu_{t}^{j})$ with respect to $1$-Wasserstein distance is equal to 1 for all $j$. Also recalling [2, Theorem 3.3], we write
$1=|\mu^{j}|_{1\textrm{-}\mathrm{var}}=\int|\gamma|_{1\textrm{-}\mathrm{var}}{\mathrm{\,d}}\pi^{j}=\int|\gamma|_{1\textrm{-}\mathrm{var}}{\mathrm{\,d}}\pi=|\mu|_{1\textrm{-}\mathrm{var}}=1.$ (4.38)
Now take $p>1$ and note that the $p$-variation of the curve $(\mu_{t}^{j})$ with respect to $p$-Wasserstein metric is achieved by the partition $(\frac{i}{2^{j+1}})_{i\in\{0,\cdots 2^{j+1}\}}$. This can be verified using the interpolation result [13, Proposition 5.5]. We thus obtain
@@
Observation 4 ($W^{1,p}$-regularity of $(\mu_{t}^{j})$ remains constant). Consider $1<p<\infty$. Each $(\mu_{t}^{j})$ is an absolutely continuous curve in $p$-Wasserstein space with its metric speed $|\dot{\mu}_{t}^{j}|=1$ for almost all time for all $j$. Recalling [17, Theorem 5], we have here:
$1=|\mu^{j}|^{p}_{W^{1,p}}=\int|\gamma|^{p}_{W^{1,p}}{\mathrm{\,d}}\pi^{j}=\int|\gamma|^{p}_{W^{1,p}}{\mathrm{\,d}}\pi=|\mu|^{p}_{W^{1,p}}=1.$ (4.40)
Observation 5 ($W^{\alpha,p}$-regularity of $(\mu_{t}^{j})$ decreases in $j$). Consider $1<p<\infty$ and $1/p<\alpha<1$. Then
@@
This is the main counterexample of the paper, whose goal is to show that if we are given $(\mu_{t})\in W^{\alpha,p}(I;P_{p}(\mathcal{X}))$ with $1/p<\alpha<1$ but without the compatibility property, there can be no lift with finite $W^{\alpha,p}$-energy. Specifically, we construct a non-compatible curve $(\mu_{t})_{t\in[0,1]}$ in a $p$-Wasserstein space with finite $W^{\alpha,p}$-norm whose only existing lift $\pi$ has infinite $W^{\alpha,p}$-energy:
@@
The construction follows by combining Examples 4.4 and 4.3, leading to an example on a cylinder. Due to the similarities, we shall not repeat all the details. Here, by choosing $\upgamma=\alpha$ for the mass $w_{j}$ in (4.13), we let the $b^{\alpha,p}$-energy (4.23) of the lift $\pi$ diverge. Meanwhile, by modifying the measures with the particle splitting procedure explained in Example 4.4, we let the $b^{\alpha,p}$-norm of $(\mu_{t})$ decrease and eventually converge to a finite quantity. This procedure clearly won't change the energy. In particular, the computations in the aforementioned examples can be used.
@@
$w_{j}\coloneqq\bar{w}\,2^{-jp\alpha},\qquad\bar{w}(p,\alpha)\coloneqq 1-2^{-p\alpha},$ (4.43)
@@
$\int_{\Gamma}|\gamma|^{p}_{W^{1,p}}{\mathrm{\,d}}\pi(\gamma)$ $=\int_{\Gamma}\int_{0}^{1}|\dot{\gamma}_{t}|^{p}{\mathrm{\,d}}t{\mathrm{\,d}}\pi(\gamma)=\bar{w}2^{p}\sum\nolimits_{j=0}^{\infty}2^{j(p-\alpha p)}=+\infty.$ (4.46)
@@
$\int_{\Gamma}|\gamma|^{p}_{b^{\alpha,p}}{\mathrm{\,d}}\pi(\gamma)=\frac{\bar{w}}{2^{-\alpha p}-2^{-p}}\sum\nolimits_{j=0}^{\infty}2^{j(\alpha p-\alpha p)}=+\infty.$ (4.47)
@@
$|\mu^{j}|^{p}_{b^{\alpha,p}}=D_{j}^{p}\,2^{n\alpha p}+D_{j}^{p}\,\frac{2^{n\alpha p}}{2^{p-\alpha p}-1}=\frac{2^{2\alpha p-p}}{1-2^{\alpha p-p}}\cdot 2^{j(2\alpha p-p)}.$ (4.49)